OpenStudy (anonymous):

at the beginning of a bobsled race, teammates push their bobsled along the track. a. before they jump in, they push the bobsled(m=150kg) from rest to a speed of 4.2 m/s. How much kinetic energy does the bobsled gain? b. how much work did the bobsled team do? Explain c. if they push the bobsled with a force of 500 N, over what distance did they apply the force? d. Later, its racing down the track at a speed of 28 m/s. with the team in it, the bobsled has a total mass of 430 kg. How much work must be done by air resistance to slow down the bobsled to 25m/s?

OpenStudy (anonymous):

a. Ek = 0.5*mv^2 m=150 v=4.2 Ek = 0.5*150*(4.2^2) = 1323 J b. I believe work (W) is equal to the change in energy (E) so the equation is just W=(delta)E, so 1323 J again? c. W=Fd is the equation, which can be rearranged to d=W/F d=1323/500 d=2.646m d. m = 430 velocity (initial) = 28 velocity (final) = 25 Ek (initial) = 0.5*m*v^2 = 0.5*430*(28^2) =168560 J Ek (final) = 0.5*m*v^2 = 0.5*430*(25^2) = 134375 J change in energy = 168560 - 134375 = 34185 J Work (W) = change in energy (E) Work = 34185 J or, 34.185 KJ beware, I might have made a mistake in that last bit.

OpenStudy (anonymous):

Thank you @youngelk